Please Help Me With this Question. "Solve" Xsquared + 5x + 1 = 0
Answer by stanbon(75887)

Xsquared + 5x + 1 = 0
---
x^2 + 5x + 1 = 0
---
Using the Quadratic Formula you get:
x = [-5 +- sqrt(25 - 5*1*1)]/2
x = [-5 +- sqrt(20)]/2
x = (-5 + 2sqrt(5))/2 or x = (-5 - 2sqrt(5)]/2
====================================================
Cheers,
Stan H.